An amount of shs 2000 is invested at an interest if 5% per month. If shs 200 is added at the beginning of each successive month but with no withdraws,
1.Give an expression for the value accumulated after n months.
2.After how many months will the amount have accumulated first exceed shs. 42000?

Answer by Boreal(15235)

Total=2000(1+(.05/12))^n-1/(0.05/12)
42000=2000(1+(.05/12))^n-1/(0.05/12)
divide by 2000 and multiply by (0.05/12)
0.0875=1.00416667^n-1
1.0875=1.00416667^n
ln both sides
0.0839=n *0.00416
n=20.18 months; 21 months ANSWER
Check with 20 and 21
$41623.63 and $43797.06